OpenStudy (tylermckinney16):

1.Explain what it means to solve a right triangle. 2.Create your own example problem. 3.Show the worked-out solutions to the problem you created.

OpenStudy (welshfella):

1 to completely solve it means to find the measures of all 3 angles and sides, given some of these measures

OpenStudy (welshfella):

2 we might be given the triangle below 

OpenStudy (welshfella):

how could we find the length of the hypotenuse x?

OpenStudy (welshfella):

what theorem would we use to do this?

OpenStudy (tylermckinney16):

i think the trig one lol

OpenStudy (welshfella):

yes you could use a trig identity but what about a well known geometrical method? - it starts with a P.

OpenStudy (tylermckinney16):

Pemdas?

OpenStudy (welshfella):

No. He's a famous Greek mathematician and philosopher.

OpenStudy (welshfella):

c^2 = a^2 + b^2

OpenStudy (welshfella):

- for a right angled triangle where c = the longest side.

OpenStudy (tylermckinney16):

ok

OpenStudy (welshfella):

Pythagoras

OpenStudy (tylermckinney16):

yeah lol i remember now.

OpenStudy (welshfella):

so can you work out what measure of x is?

OpenStudy (tylermckinney16):

i can try yeah hold on.

OpenStudy (tylermckinney16):

is it 1?

OpenStudy (welshfella):

i don't think so - the longest side can't be 1 if the other 2 are 3 and 4

OpenStudy (tylermckinney16):

5?

OpenStudy (welshfella):

x^2 = 3^2 + 4^2 find x.

OpenStudy (tylermckinney16):

oh ok i got it hold on

OpenStudy (tylermckinney16):

25

OpenStudy (welshfella):

x^2 = 25 so x = square root 25

OpenStudy (tylermckinney16):

So is that are answer?

OpenStudy (welshfella):

thats the value of the other side. Now to find the angles you can use one of the trig ratios sin, cos or tan.

OpenStudy (tylermckinney16):

ok

OpenStudy (welshfella):